An alcohol solution was made by mixing 12 fl. oz of a 30% alcohol solution and 3 fl.oz of a 90% solution. What is the concentration of the mixture?
Answer by TimothyLamb(4379)

---
30(12) + 90(3) = x(15)
360 + 270 = x(15)
630 = x(15)
---
answer:
x = 42% alcohol in the final mixture
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php